# TabEBM: A Tabular Data Augmentation Method with Distinct Class-Specific Energy-Based Models

### v1

@@ 4 Discussion & Related Work
Generative Models for Tabular Data. The common paradigm for tabular data generation is to adapt Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) [86, 63]. For instance, TableGAN employs a convolutional neural network to optimise the label quality [63], and TVAE is introduced in [86] as a variant of VAE for tabular data. However, these methods learn the joint distribution and thus cannot preserve the stratification of the original data (Appendix B). CTGAN [86] refines the generation to be class-conditional. The recent ARF [84] is an adversarial variant of random forest for density estimation, and GOGGLE [47] enhances VAE by learning relational structure with a Graph Neural Network (GNN). Some recent work focuses on generation with denoising diffusion models [42, 88, 40, 44]. For instance, TabDDPM [42] demonstrates that diffusion models can approximate typical distributions of tabular data. Although these class-conditional models can preserve the label distribution, they struggle to outperform Baseline and standard SMOTE in data augmentation [72, 48].
We attribute the performance degradation in current class-conditional models to their reliance on a single shared model to approximate all class-conditional densities. For instance, another promising generative approach uses pre-trained models like Prior-Data Fitted Networks (PFNs), and the recent TabPFGen [48] adapts such models into one shared class-conditional generator. However, TabPFGen’s shared generator can lead to inaccurate density estimates, particularly in high-noise and class-imbalance situations (see examples in Appendix B). As noise increases, TabPFGen’s inferred densities fluctuate significantly and diverge from the true data distributions. In contrast, TabEBM uses class-specific EBMs to model each class’s marginal distributions, and the results in Appendix B reveal that our design choice reduces the impact of noise and data imbalance. TabEBM focuses on approximating and generating for one class at a time, remaining unaffected by noise from other classes. Overall, our results demonstrate that TabEBM consistently improves performance across different datasets and sample sizes, outperforming TabPFGen. Moreover, TabPFGen is limited in usability (e.g., it supports only up to ten classes), while TabEBM scales to any number of classes.
@@
Limitations and Future Work. TabEBM is a general method that relies on an underlying binary classifier, and as such, its strengths and weaknesses are directly tied to this classifier. We used TabPFN because it is a well-established open-source pre-trained model for tabular data. Therefore, TabEBM inherits some of TabPFN’s limitations, particularly in scaling to a larger number of features. TabEBM can handle datasets with over 1000 samples, overcoming TabPFN’s limitation, as it processes one class at a time. In Section C.5.3, we show that TabEBM outperforms other generators on larger datasets, though the performance gains decrease as the sample size increases. Although TabPFN has a 10-class limit, TabEBM can handle an unlimited number of classes by training TabPFN on binary surrogate tasks instead of the original multi-class problem. Our results also show that TabEBM can handle categorical data by encoding categorical features with leave-one-out target statistics. We stress that TabEBM is compatible with any classifier that can be adapted into EBMs, as described in Section 2. As foundational models for tabular data evolve [82], new models capable of handling more features and samples are expected. Integrating them into TabEBM will enhance its ability to manage high-dimensional datasets, increasing its versatility and utility.
@@ 5 Conclusion
Broader Impact Statement
@@
Appendix A Reproducibility
A.1 Datasets
@@ A.3 Data Preprocessing
A.4 Software and Computing Resources
@@
We ensure the consistency and reproducibility of experimental results by implementing a uniform pipeline using PyTorch Lightning, an open-source library under an Apache-2.0 licence. We further fixed the random seeds for data loading and evaluation throughout the training and evaluation process. This ensured that TabEBM and all benchmark models were trained and evaluated on the same set of samples. We also attach our code to this submission and will release it under the MIT licence upon publication. The experimental environment settings, including library dependencies, are specified in the associated code for reference and reproduction purposes.
@@
A.5 TabEBM open-source library
We implemented TabEBM as an extensible library, included as a zip file in the Supplementary material. The library will be made publicly available after publication. For practitioners, it offers an easy-to-use, domain-agnostic tool that requires no training, making it particularly suitable for data augmentation, especially in small datasets. For researchers, the library includes the complete implementation of TabEBM, facilitating future extensions and investigations into class-specific energy-based models.
@@ A.5 TabEBM open-source library
    from TabEBM import TabEBM
@@
    Compute and visualise the energy function: The library allows computation of TabEBM’s energy function and the unnormalised data density. The demo notebook, TabEBM_approximated_density.ipynb, shows the TabEBM-inferred densities under conditions of data noise and class imbalance (thus recreating the plots from Appendix B).
A.6 Implementation of Generators
TabEBM. In all our experiments, the surrogate binary classifier in TabEBM is a pretrained in-context model, TabPFN [33], using the official model weights released by the authors (https://github.com/automl/TabPFN/raw/main/tabpfn/models_diff/prior_diff_real_checkpoint_n_0_epoch_42.cpkt). We use TabPFN with three ensembles. We use four surrogate negative samples, $\mathcal{X}_{c}^{\text{neg}}$, positioned at $\alpha^{\text{neg}}_{\text{dist}}=5$ standard deviations from zero, in random corners of a hypercube in ${\mathbb{R}}^{D}$ (as explained in Section 2.2), distant from any real data. In Section C.1, we show that TabEBM is robust to the distribution of the negative samples.
We use SGLD [85] for sampling from TabEBM, where the starting points ${\mathbf{x}}_{0}^{\text{synth}}$ are initialised by adding Gaussian noise with zero mean and standard deviation $\sigma_{\text{start}}=0.01$ to a randomly selected sample of the specific class, i.e., ${\mathbf{x}}_{0}^{\text{synth}}\sim\mathcal{N}(\mathcal{X}_{c},\sigma_{\text{start}}^{2}\mathbf{I})$. For SGLD, we used the following parameters: step size $\alpha_{\text{step}}=0.1$, noise scale $\alpha_{\text{noise}}=0.01$ and number of steps $T=200$. We found TabEBM to be robust to the SGLD settings (see Section C.2). We attach the implementation code for TabEBM and will release it as a public open-source library available after publication (see Section A.5 for details on our library).
@@ A.6 Implementation of Generators
A.7 Implementation of Downstream Predictors
@@
Appendix B Limitations of Existing Generative Methods
@@
Appendix C Extended Experimental Results
C.1 Ablations on the distribution of the surrogate negative samples
C.1.1 Ablations on placing the negative samples
@@
C.1.2 Varying the number of negative samples
@@
C.1.3 Varying the distance of the negative samples
@@
C.2 Ablations on the sensitivity to the hyperparameters of SGLD sampling
@@
C.3 Distribution of Logits and Unnormalized Density in TabEBM
@@
C.4 Complete Trade-off Figures with Error Bars
@@
C.5 Results on Data Augmentation
C.5.1 Results on eight OpenML datasets.
@@ C.5.2 Results on six UCI Datasets
C.5.3 Results on larger sample sizes
@@
C.6 Results on Statistical Fidelity
@@
C.6.1 Similarity between Real Train Data and Synthetic Data
@@
C.6.2 Similarity between Real Test Data and Synthetic Data
@@
C.7 Results on Privacy Preservation
C.7.1 Downstream Accuracy in Data Sharing
@@ C.7.2 DCR Evaluation
C.7.3 Delta-presence Evaluation

### v2

@@ 4 Discussion & Related Work
Generative Models for Tabular Data. The common paradigm for tabular data generation is to adapt Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) [85, 63]. For instance, TableGAN employs a convolutional neural network to optimise the label quality [63], and TVAE is introduced in [85] as a variant of VAE for tabular data. However, these methods learn the joint distribution and thus cannot preserve the stratification of the original data (Appendix C). CTGAN [85] refines the generation to be class-conditional. The recent ARF [83] is an adversarial variant of random forest for density estimation, and GOGGLE [47] enhances VAE by learning relational structure with a Graph Neural Network (GNN). Some recent work focuses on generation with denoising diffusion models [42, 87, 40, 44]. For instance, TabDDPM [42] demonstrates that diffusion models can approximate typical distributions of tabular data. Although these class-conditional models can preserve the label distribution, they struggle to outperform Baseline and standard SMOTE in data augmentation [71, 48].
We attribute the performance degradation in current class-conditional models to their reliance on a single shared model to approximate all class-conditional densities. For instance, another promising generative approach uses pre-trained models like Prior-Data Fitted Networks (PFNs), and the recent TabPFGen [48] adapts such models into one shared class-conditional generator. However, TabPFGen’s shared generator can lead to inaccurate density estimates, particularly in high-noise and class-imbalance situations (see examples in Appendix C). As noise increases, TabPFGen’s inferred densities fluctuate significantly and diverge from the true data distributions. In contrast, TabEBM uses class-specific EBMs to model each class’s marginal distributions, and the results in Appendix C reveal that our design choice reduces the impact of noise and data imbalance. TabEBM focuses on approximating and generating for one class at a time, remaining unaffected by noise from other classes. Overall, our results demonstrate that TabEBM consistently improves performance across different datasets and sample sizes, outperforming TabPFGen. Moreover, TabPFGen is limited in usability (e.g., it supports only up to ten classes), while TabEBM scales to any number of classes.
@@
Limitations and Future Work. TabEBM is a general method that relies on an underlying binary classifier, and as such, its strengths and weaknesses are directly tied to this classifier. We used TabPFN because it is a well-established open-source pre-trained model for tabular data. Therefore, TabEBM inherits some of TabPFN’s limitations, particularly in scaling to a larger number of features. TabEBM can handle datasets with over 1000 samples, overcoming TabPFN’s limitation, as it processes one class at a time. In Section D.5.3, we show that TabEBM outperforms other generators on larger datasets, though the performance gains decrease as the sample size increases. Although we implement TabEBM with TabPFN in this paper, we stress that TabEBM is compatible with any classifier that can be adapted into EBMs, as described in Section 2. As foundational models for tabular data evolve [81], new models capable of handling more features and samples are expected. Integrating them into TabEBM will enhance its ability to manage high-dimensional datasets, increasing its versatility and utility. In addition, we would like to point out the imperfection of widely adopted univariate statistical fidelity metrics (i.e., Inverse KL, KS test and $\chi^{2}$ test). However, evaluating generators’ ability to capture the joining feature relationships remains an open research question [78]. We leave this for future work to explore.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
The authors would like to thank Francisco Vargas, Randall Balestriero, and Otilia Stretcu for their insightful discussions and valuable input early in the project. NS acknowledges the support of the U.S. Army Medical Research and Development Command of the Department of Defense; through the FY22 Breast Cancer Research Program of the Congressionally Directed Medical Research Programs, Clinical Research Extension Award GRANT13769713. Opinions, interpretations, conclusions, and recommendations are those of the authors and are not necessarily endorsed by the Department of Defense.
@@
Appendix A Broader Impact Statement
Appendix B Reproducibility
B.1 Datasets
@@ B.3 Data Preprocessing
B.4 Software and Computing Resources
@@
We ensure the consistency and reproducibility of experimental results by implementing a uniform pipeline using PyTorch Lightning, an open-source library under an Apache-2.0 licence. We further fixed the random seeds for data loading and evaluation throughout the training and evaluation process. This ensured that TabEBM and all benchmark models were trained and evaluated on the same set of samples. The experimental environment settings, including library dependencies, are specified in the open-source library for reference and reproduction purposes.
@@
B.5 TabEBM open-source library
We implemented TabEBM as an extensible library, and the code is available on https://github.com/andreimargeloiu/TabEBM. For practitioners, it offers an easy-to-use, domain-agnostic tool that requires no training, making it particularly suitable for data augmentation, especially in small datasets. For researchers, the library includes the complete implementation of TabEBM, facilitating future extensions and investigations into class-specific energy-based models.
@@ B.5 TabEBM open-source library
    from tabebm.TabEBM import TabEBM
@@
    Compute and visualise the energy function: The library allows computation of TabEBM’s energy function and the unnormalised data density. The demo notebook, TabEBM_approximated_density.ipynb, shows the TabEBM-inferred densities under conditions of data noise and class imbalance (thus recreating the plots from Appendix C).
B.6 Implementation of Generators
TabEBM. In all our experiments, the surrogate binary classifier in TabEBM is a pretrained in-context model, TabPFN [33], using the official model weights released by the authors (https://github.com/automl/TabPFN/raw/main/tabpfn/models_diff/prior_diff_real_checkpoint_n_0_epoch_42.cpkt). We use TabPFN with three ensembles. We use four surrogate negative samples, $\mathcal{X}_{c}^{\text{neg}}$, positioned at $\alpha^{\text{neg}}_{\text{dist}}=5$ standard deviations from zero, in random corners of a hypercube in ${\mathbb{R}}^{D}$ (as explained in Section 2.2), distant from any real data. In Section D.1, we show that TabEBM is robust to the distribution of the negative samples.
We use SGLD [84] for sampling from TabEBM, where the starting points ${\mathbf{x}}_{0}^{\text{synth}}$ are initialised by adding Gaussian noise with zero mean and standard deviation $\sigma_{\text{start}}=0.01$ to a randomly selected sample of the specific class, i.e., ${\mathbf{x}}_{0}^{\text{synth}}\sim\mathcal{N}(\mathcal{X}_{c},\sigma_{\text{start}}^{2}\mathbf{I})$. For SGLD, we used the following parameters: step size $\alpha_{\text{step}}=0.1$, noise scale $\alpha_{\text{noise}}=0.01$ and number of steps $T=200$. We found TabEBM to be robust to the SGLD settings (see Section D.2).
@@ B.6 Implementation of Generators
B.7 Implementation of Downstream Predictors
@@
Appendix C Limitations of Existing Generative Methods
@@
Appendix D Extended Experimental Results
D.1 Ablations on the distribution of the surrogate negative samples
D.1.1 Ablations on placing the negative samples
@@
D.1.2 Varying the number of negative samples
@@
D.1.3 Varying the distance of the negative samples
@@
D.2 Ablations on the sensitivity to the hyperparameters of SGLD sampling
@@
D.3 Distribution of Logits and Unnormalized Density in TabEBM
@@
D.4 Complete Trade-off Figures with Error Bars
@@
D.5 Results on Data Augmentation
D.5.1 Results on eight OpenML datasets.
@@ D.5.2 Results on six UCI Datasets
D.5.3 Results on larger sample sizes
@@
D.6 Results on Statistical Fidelity
@@
D.6.1 Similarity between Real Train Data and Synthetic Data
@@
D.6.2 Similarity between Real Test Data and Synthetic Data
@@
D.7 Results on Privacy Preservation
D.7.1 Downstream Accuracy in Data Sharing
@@ D.7.2 DCR Evaluation
D.7.3 Delta-presence Evaluation
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Please refer to Section 1, where we state our research topic and objective and then connect them to our contributions. We also provide an open-source library of the proposed method (https://github.com/andreimargeloiu/TabEBM).
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Please refer to Section 4, where we analyse the limitations of the proposed model and promising directions for future research.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: Please refer to Section 2, where we provide full theoretical derivations of our proposed method.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: Please refer to Appendix B, where we provide full details on reproducing the results in the paper. Furthermore, we have also provided the open-source implementation of our proposed method.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: Please refer to Appendix B. All datasets used in this paper are publicly available, and the implementations of benchmark generators are open-source. In addition, we have also provided the implementation of our proposed method.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Please refer to Appendix B, where we provide full descriptions of the experimental setup to reproduce the results in the paper.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: Please refer to Section 3, where we provide standard deviations for all tables. Figure 6 and Figure 4 (Right) contain error bars. Due to the page limit, the error bars for all other figures are available in Appendix D. In Section 3.2 and Section D.6, we show statistical significance tests of the similarity between real data and synthetic data.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: Please refer to Section B.4, where we provide full details on the computation resources used in the paper.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We carefully check the NeurIPS Code of Ethics, and we confirm that our work follows the Code in every respect.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: Please refer to Appendix A, where we include the societal impacts of our work.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: [N/A]
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: Please refer to Appendix B, where we provide the open-source licenses followed by the creators or original owners of assets.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We provide the implementation of our method as a python library attached to this submission. We will make it publicly available post-publication.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: [N/A]
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [No]
  60. Justification:[N/A]